Mae purchased a home for ₱4,570,000 and obtained a 20-year loan with an annual interest rate of 8.5%, after paying a down payment of 25%. Find the the amount of interest paid on the loan over the 20 years

1
Expert's answer
2022-02-01T17:00:58-0500

"P\ =\frac{A\left(\frac{r}{n}\right)}{{1\ -\ \left(1\ +\ \frac{r}{n}\right)}^{-nt}}"


"where\ A\ =\ \frac{25}{100}\ \times\ 4,570,000\ =\ 1,142,500"


"where\ r\ =\ \frac{8.5}{100}\ =\ 0.085"


"t\ =\ 20years"


"n\ =\ 12\ months"


"P\ =\frac{1,142,500\left(\frac{0.085}{12}\right)}{{1\ -\ \left(1\ +\ \frac{0.085}{12}\right)}^{-12\times20}}"


"P\ =\frac{8,092.708}{1\ -\ \left(0.1838\right)}"


"P\ =\ 9,915.10"


"The\ amount\ of\ interest\ paid\ on\ the\ loan\ over\ 20years\ will\ be\ \ ₱9,915.10"